A man made a motor trip of 220 mile in 5 hours. Part of the time he averaged 30 mph and part of the time he averaged 50 mph. How many hours did he travel at each rate
Answer by Paul(988)

Total duration ar 30mph = 30(5-x)
Total duration at 50mph = 50x
50x+30(5-x)=220
50x-30x+150=220
20x=70
x=3.5

Total duration at 50mph = 50(5-x)
Total duration at 30mph = 30x
30x+50(5-x)=220
30x+250-50x=220
-20x=-30
x=1.5

Hence for 30mph the duration was for 3.5h and for 50mph the duration was for about 1.5h
Paul.